### v2

@@ II-C2 Statistical analysis
These values are calculated for each pulse of a signal. So in our case, the variable X represents a vector of all the samples in a pulse. If the shape of the cycle changes, then these statistical values will no longer be constant. To be able to detect outliers, thresholds that detect skewness, kurtosis and standard deviation values that are not normal, i.e. values for artifact-free cycles, were set up. For this reason, the distribution of each of these three statistics over a pulse can be estimated using a normal distribution [28]. The aim is to reduce the risk of a pulse being incorrectly annotated. To do this, a wide confidence interval is taken to ensure that the probability that the value of the corresponding statistic is not unnecessarily rejected. If X is considered to be a variable which can be approximated by a normal distribution $\mathcal{N}(\mu,\,\sigma^{2})$, the probability that this variable lies within the chosen confidence interval can be written as follows:
@@
After several experiments, this 95% confidence interval gives the best results, as it reduces the risk of poor detection. So a lower and an upper thresholds can be defined as follows:
@@
With the mean and standard deviation calculated for each statistic measure by taking the set of values for each pulse of a signal. To effectively detect motion artifacts, a waveform segment is classified as containing motion artifacts if at least one of the three statistics falls outside the defined thresholds. The result of this first step is a small proportion of the annotated dataset, with a binary value for each pulse: pulse with artifact or without artifact. The annotations given by the algorithm are then compared with the expert’s annotations and found to have 80% similarity. After examining the annotations with the expert, the function chosen to segment the pulses did not always correctly segment a pulse that was formed by a distinct diastole and systole curve. In a cardiac cycle, diastole is the relaxation phase when the heart fills with blood, and systole is the contraction phase when the heart pumps blood out to the body or lungs. In this case, two pulses were detected instead of one. This segmentation error partly explains the 20% difference in annotation between the expert and the algorithm. The percentage of similarity is considered high enough to validate the expert’s annotations.
II-C3 Imbalanced dataset
@@
The two classes of annotated data are unevenly distributed. In fact, the annotation includes many more pulses without motion artifacts, approximately 80% and 20% of pulses with motion artifacts. For accurate results with the algorithms, the data needs to be resampled. The complex characteristics of our clinical data: small training size, many features, and correlations between the features, makes the task more complicated. Understanding the interconnectedness of these variables is crucial for accurate analysis and prediction [17]. Oversampling and undersampling methods are the most frequently used. Under-sampling reduces the majority class examples, achieving a balanced dataset, with random under-sampling (RUS) being a well-known method. However, under-sampling may lead to the loss of valuable information from the majority class. On the other hand, over-sampling increases the minority class examples. Random over-sampling (ROS) replicates existing minority examples, but it may result in overfitting. Synthetic minority over-sampling technique (SMOTE) generates artificial minority examples by interpolating between selected examples and their nearest neighbors. Modifications such as adaptive synthetic sampling (ADASYN) adjust the number of artificial minority examples based on the density of majority examples surrounding the original minority example [29]. Also, it is concluded that there is no clear winner between oversampling and undersampling to compensate for the class imbalance if factors such as class distribution, class prevalence, and features correlations in medical decision-making [17] are not taken into consideration. In the section IV the different results obtained with the sampling methods will be presented, to conclude on the best method for our study.
II-C4 Label Propagation
@@
The Label Propagation algorithm is an iterative algorithm that assigns labels to unlabeled data points by propagating labels through the dataset [30]. In graph-based semi-supervised learning methods, a graph where each node is represented by a vector of features is created. The edges between nodes are weighted based on how similar the features are. When the weights of the edges are high, it means that the connected nodes are likely to have the same label. This idea is based on the assumption that samples close to each other in the graph are part of the same group or category [31]. At the start of the algorithm, only a small proportion of the data is already labeled, corresponding here to the proportion of data annotated in the previous step. In our case, considering that we have around 51 pulses per signal and that we have annotated 10% of the entire database of 1571 signals, we therefore have 8000 pulses, and thus 8000 nodes in the graph. This algorithm is based on the hypothesis that if two nodes are connected that means they carry a similarity. Usually, the Euclidean distance between nodes is calculated to establish the graph. Depending on the kernels chosen for the algorithm’s operation, this distance measurement may be different. Consider the following notations:
@@
The aim of this algorithm is to look, in the final state, at all the probabilities a node has of belonging to a certain class and to take the largest. $Y$ a matrix with rows containing the probabilities that a node belongs to a certain class is considered. This matrix $Y$ is a $N\times k$ matrix where $N=l+u$. Also considered $T$, a $N\times N$ probability transition matrix. This matrix T is obtained by calculating the degree matrix ($D$) and the adjacency matrix ($A$). It defines the probability of jumping from one node to another in $t$ steps. This number $t$ can tend towards infinity [32]. The matrix $Y$ contains two sub-matrices: $Y_{l}$ and $Y_{u}$, respectively for the known and unknown labels. The same applies to the $T$ matrix, which contains 4 sub-matrices:
  • $T_{ll}$: probability to get from labelled nodes to labelled nodes. This matrix will be an identity matrix.
  • $T_{lu}$: probability to get from labelled nodes to unlabelled nodes. This will be a zero matrix because labelled nodes are absorbing states, it means you are in a self-loop and can’t move in any direction.
  • $T_{ul}$ and $T_{uu}$: probability to get from unlabelled nodes to labelled and unlabelled nodes respectively.
Consider $\hat{Y}$, the probability matrix of annotations obtained in the final state. The matrix T is set to the infinite power and $Y_{0}$ represents the initial annotations of the nodes. The equation for the final state of this algorithm can be expressed as:
@@
In 7, the matrix $T$ is set to the power $t$ with $t$ tending to infinity. It can be written:
@@
The sum between the brackets is similar, when $t$ tends to infinity, to a geometric series that has an argument that is less than 1 in modulus. And if $T_{uu}$ is multiplied by itself a large number of times, knowing that the values are less than 1, it will become very close to 0. Therefore, a conclusion on the limit of the transition matrix for a very large number of steps, is:
@@
    Calculation of the probability transition matrix $T$. This matrix is linked to the degree matrix $D$, it is a diagonal matrix where each diagonal element corresponds to the sum of edge weights connected to that node. And also linked to the adjacency matrix $A$, it is a square matrix where each row and column corresponds to a node, and the value at the intersection indicates whether there’s an edge (value 1, otherwise 0) connecting those nodes. The formula is: $T=D^{-1}\cdot A$. This matrix is the same throughout the algorithm.
@@
While performing label propagation, groups of closely linked nodes quickly reach a consensus on a single label, causing many labels to vanish. This leaves only a few labels remaining after propagation. When nodes end up with the same label after convergence, it signifies that they are part of the same group.
II-D Classification
Once the ground truth has been established, the aim is to classify the pulses and compare the results obtained with the annotations. Machine learning classifiers are used for classification. These automatic algorithms categorize data into the two classes of our problem. They operate as mathematical models, utilizing statistical analysis and optimization techniques to detect patterns within the data. By identifying these patterns, classifiers can assign each instance to a specific class or category. There are a wide variety of traditional classifiers, both supervised and unsupervised. To process medical data, which is also temporal data, supervised classifiers have been chosen to be utilized. Here are 4 examples [21]:
@@ II-D Classification
    K-Nearest Neighbor (KNN): this is a supervised method where k represents the number of neighbors. For classification, when given a new input data point, the algorithm identifies the k nearest neighbors from the training dataset based on their feature similarity. The class label of the majority of these k neighbors is then assigned to the new data point.
@@
    Support Vector Machine (SVM): in SVM, data points are mapped as vectors within a high-dimensional space. The algorithm aims to identify the optimal hyperplane that distinctly categorizes the classes. In binary classification, a hyperplane can be considered as a boundary delineating two distinct data classes. While numerous hyperplanes might achieve this separation, the algorithm selects the one that provides the most effective separation. For a specific classification purpose, as is the case for this project, we have subsequently used SVC (Support Vector Classification) which is a type of SVM specialized for classifications.
@@
    Decision Tree: Each internal node represents a feature or attribute, and each branch represents a decision rule based on that feature. The leaf nodes of the tree represent the final class label or predicted value. During training, each value is separated based on the attribute. When making predictions, new data points traverse the decision tree by following the decision rules at each node until reaching a leaf node, which then provides the predicted class label or value.
@@
    Naive Bayes classifier: this classifier use probability to predict whether an input will fit into a certain category. It builds a statistical model based on these probabilities. Naive Bayes calculates the likelihood of the data point belonging to each class using the previously estimated probabilities.
@@
Traditional classifiers have big advantages for small or medium datasets that require simpler or linear models. They have few layers in their architecture, on the other hand, deep learning (MLP and Transformers) architectures are comprised of multiple layers of neural networks. Deep architectures take advantage of unsupervised pre-training at layer level, which facilitates efficient tuning of the deep networks and enables them to extract intricate structures from input data. These extracted features at higher levels contribute to improved predictions and overall performance [33]. For classification, MLP and Transformers are neural networks classifiers:
@@
  • Multilayer Perceptron (MLP): it consists of multiple layers of nodes (neurons) that are interconnected through weighted connections. MLP employs a feedforward mechanism, where information flows from the input layer through the hidden layers to the output layer. Each node in the network applies an activation function to the weighted sum of its inputs to produce an output. Through a process called backpropagation, the MLP classifier adjusts the weights to minimize the error between predicted and actual labels during training.
@@
  • Transformers: it relies on the attention mechanism. The attention-mechanism looks at an input sequence and decides at each step which other parts of the sequence are important. Transformer is an architecture for transforming one sequence into another one with the help of two parts (Encoder and Decoder).
@@
The objective is to apply all these classifiers to the PPG signal pulses so that a comparison of the classifiers on our medical data can be built. In addition to being compared with each other, these classifiers will also be compared with the Label Propagation semi-supervised algorithm, which not only annotates our database but also classifies artifacts in PPG signals.
III Experimental Implementation
@@
First, as a reminder, in the Label Propagation algorithm, the two input matrices are the annotation matrix, which is a binary vector, and a matrix containing the features for each pulse. Each pulse represents a node in the algorithm’s graph. For the choice of features, the signal from a temporal perspective has been considered. Therefore, an input matrix for the algorithm of size 256 samples $\times$ the number of pulses, can be obtained.
@@
To be able to evaluate our results different metrics have been chosen. The negative state (0) is a pulse without motion artifacts, whereas the positive state (1) is a pulse with motion artifacts. All these measures are based on the evaluation of false negatives (FN), pulse with artifact incorrectly identified as a clean pulse, false positives (FP), clean pulse incorrectly identified as a pulse with artifact, true negatives (TN), clean pulse correctly identified as a clean pulse, and true positives (TP), pulse with artifact correctly identified as a pulse with artifact. The following metrics are defined:
@@
  • Precision, Recall & F1: these three scores give a more general idea of how the algorithm works, rather than just looking at the algorithm’s accuracy, which can be biased in certain situations. They are defined as follows:
@@
  • AUROC: the AUROC (Area Under the Receiver Operating Characteristic curve) represents the probability that the model correctly ranks a randomly chosen positive instance higher than a randomly chosen negative instance. The ROC curve is created by plotting the TP rate against the FP rate.
@@
Using these metrics, the hyperparameters of our model for the Label Propagation algorithm need to be defined. The first parameters defined are the parameters of the function used. The choice of kernel is between KNN or RBF (radial basis function), and depending on this choice, two other associated parameters could be modified. The maximum number of iterations and the algorithm’s convergence tolerance remained the default values: 1000 iterations and $10^{-3}$ for the threshold of convergence. The first parameter to be defined was the choice of kernel. For this, a cross-validation was carried out on the data. This involves dividing the data into several parts, then running the two algorithms using different values for the parameters on each part, keeping one part aside for performance testing. Then a calculation of the average performance can be done over all the test parts for each value and choose the one that gives the best performance. A KNN kernel with a number of neighbors of 7 has been chosen. The data are separated as follows: 70% training and 30% testing. The data from the training part are redivided evenly to obtain 50% of unlabeled data and 50 % of labeled data.
IV Results and Discussion

### v3

@@ III-C2 Statistical analysis
These values are calculated for each pulse of a signal. So, in our case, the variable X represents a vector of all the samples in a pulse. If the shape of the cycle changes, then these statistical values will no longer be constant. To be able to detect outliers, thresholds that detect skewness, kurtosis, and standard deviation values that are not normal, i.e. values for artifact-free cycles, were set up. For this reason, the distribution of each of these three statistics over a pulse can be estimated using a normal distribution [31]. The aim is to reduce the risk of a pulse being incorrectly annotated. To do this, a wide confidence interval is taken to ensure that the probability that the value of the corresponding statistic is not unnecessarily rejected. If X is considered to be a variable that can be approximated by a normal distribution $\mathcal{N}(\mu,\,\sigma^{2})$, the probability that this variable lies within the chosen confidence interval can be written as follows:
@@
After several experiments, this 95% confidence interval gives the best results, as it reduces the risk of poor detection. So, lower and upper thresholds can be defined as follows:
@@
The mean and standard deviation are calculated for each statistic measured by taking the set of values for each pulse of a signal. A waveform segment is classified as containing motion artifacts to effectively detect motion artifacts if at least one of the three statistics falls outside the defined thresholds. The result of this first step is a small proportion of the annotated dataset, with a binary value for each pulse: pulse with artifact or without artifact. The annotations given by the algorithm are then compared with the expert’s annotations and found to have 80% similarity. After examining the annotations with the expert, the function chosen to segment the pulses did not always correctly segment a pulse that was formed by a distinct diastole and systole curve. In a cardiac cycle, diastole is the relaxation phase when the heart fills with blood, and systole is the contraction phase when the heart pumps blood out to the body or lungs. In this case, two pulses were detected instead of one. This segmentation error partly explains the 20% difference in annotation between the expert and the algorithm. The percentage of similarity is considered high enough to validate the expert’s annotations.
III-C3 Imbalanced dataset
@@
The two classes of annotated data are unevenly distributed. The annotation includes many more pulses without motion artifacts, approximately 80% and 20% of pulses with motion artifacts. For accurate results with the algorithms, the data needs to be resampled. The complex characteristics of our clinical data, such as small training sizes, many features, and correlations between the features, make the task more complicated. Understanding the interconnectedness of these variables is crucial for accurate analysis and prediction [23]. Oversampling and undersampling methods are the most frequently used. Under-sampling reduces the majority of class examples, achieving a balanced dataset, with random under-sampling (RUS) being a well-known method. However, under-sampling may lead to the loss of valuable information from the majority class. On the other hand, over-sampling increases the minority class examples. Random over-sampling (ROS) replicates existing minority examples, but it may result in overfitting. Synthetic minority over-sampling technique (SMOTE) generates artificial minority examples by interpolating between selected examples and their nearest neighbors. Modifications such as adaptive synthetic sampling (ADASYN) adjust the number of artificial minority examples based on the density of majority examples surrounding the original minority example [32]. Also, it is concluded that there is no clear winner between oversampling and undersampling to compensate for the class imbalance if factors such as class distribution, class prevalence, and features correlations in medical decision-making [23] are not taken into consideration. In the section V, the different results obtained with the sampling methods will be presented to conclude on the best method for our study.
III-C4 Label Propagation
@@
The LP algorithm is an iterative algorithm that assigns labels to unlabeled data points by propagating labels through the dataset. It was first presented in an article published in 2002 by X. Zhu and Z. Ghahramani, entitled ”Learning from labeled and unlabeled data with label propagation” [33]. In graph-based semi-supervised learning methods, a graph where each node is represented by a vector of features is created. The edges between nodes are weighted based on how similar the features are. When the weights of the edges are high, it means that the connected nodes are likely to have the same label. This idea is based on the assumption that samples close to each other in the graph are part of the same group or category [34]. At the start of the algorithm, only a small proportion of the data is already labeled, corresponding here to the proportion of data annotated in the previous step. In our case, considering that we have around 51 pulses per signal and that we have annotated 10% of the entire database of 1571 signals, we therefore have 8000 pulses, and thus 8000 nodes in the graph. This algorithm is based on the hypothesis that if two nodes are connected, they carry a similarity. Usually, the Euclidean distance between nodes is calculated to establish the graph. Depending on the kernels chosen for the algorithm’s operation, this distance measurement may be different. Consider the following notations:
@@
In the final state, this algorithm aims to look at all the probabilities a node has of belonging to a certain class and take the largest. $Y$ a matrix with rows containing the probabilities that a node belongs to a certain class is considered. This matrix $Y$ is a $N\times k$ matrix where $N=l+u$. Also considered $T$, a $N\times N$ probability transition matrix. This matrix T is obtained by calculating the degree matrix ($D$) and the adjacency matrix ($A$). It defines the probability of jumping from one node to another in $t$ steps. This number $t$ can tend towards infinity [35]. The matrix $Y$ contains two sub-matrices: $Y_{l}$ and $Y_{u}$, respectively, for the known and unknown labels. The same applies to the $T$ matrix, which contains 4 sub-matrices:
  • $T_{ll}$: probability to get from labeled nodes to labeled nodes. This matrix will be an identity matrix.
  • $T_{lu}$: probability of getting from labeled to unlabelled nodes. This will be a zero matrix because labelled nodes are absorbing states, it means you are in a self-loop and can’t move in any direction.
  • $T_{ul}$ and $T_{uu}$: probability to get from unlabelled nodes to labeled and unlabelled nodes, respectively.
Consider $\hat{Y}$, the probability matrix of annotations obtained in the final state. The matrix T is set to the infinite power, and $Y_{0}$ represents the initial annotations of the nodes. The equation for the final stage of this algorithm can be expressed as:
@@
In 7, the matrix $T$ is set to the power $t$ with $t$ tending to infinity. It can be written as:
@@
The sum between the brackets is similar, when $t$ tends to infinity, to a geometric series that has an argument that is less than 1 in modulus. And if $T_{uu}$ is multiplied by itself a large number of times, knowing that the values are less than 1, it will become very close to 0. Therefore, a conclusion on the limit of the transition matrix for a very large number of steps is:
@@
    Calculation of the probability transition matrix $T$. This matrix is linked to the degree matrix $D$, a diagonal matrix where each diagonal element corresponds to the sum of edge weights connected to that node. Also linked to the adjacency matrix $A$, it is a square matrix where each row and column corresponds to a node, and the value at the intersection indicates whether there’s an edge (value 1, otherwise 0) connecting those nodes. The formula is: $T=D^{-1}\cdot A$. This matrix is the same throughout the algorithm.
@@
A concrete example of how the LP algorithm works is shown in figure 4. This example is based on a sample of synthetic data where each of the three classes is represented in a band. The KNN algorithm is unaware of the band structure and fails to propagate labels efficiently. The LP model, on the other hand, recognizes this structure and uses it to its advantage to group labels.
Figure 4: Comparison of label propagation between KNN and LP model with the ”3 Bands dataset”. (a) 3 initial annotated points (3 classes represented in green, red, and blue) and 178 non-annotated points (b) annotated dataset with KNN (c) with LP. From [33].
While performing label propagation, groups of closely linked nodes quickly reach a consensus on a single label, causing many labels to vanish. Only a few labels remain after propagation. When nodes end up with the same label after convergence, it signifies that they are part of the same group.
III-D Classification
Once the ground truth has been established, the aim is to classify the pulses and compare the results obtained with the annotations. Machine learning classifiers are used for classification. These automatic algorithms categorize data into the two classes of our problem. They operate as mathematical models, utilizing statistical analysis and optimization techniques to detect patterns within the data. By identifying these patterns, classifiers can assign each instance to a specific class or category. There are a wide variety of traditional classifiers, both supervised and unsupervised. Supervised classifiers have been chosen to be utilized to process medical data, which is also temporal data. Here are 4 examples [25]:
@@ III-D Classification
    KNN: this is a supervised method where k represents the number of neighbors. For classification, when given a new input data point, the algorithm identifies the k nearest neighbors from the training dataset based on their feature similarity. The class label of the majority of these k neighbors is then assigned to the new data point.
@@
    SVM: in SVM, data points are mapped as vectors within a high-dimensional space. The algorithm aims to identify the optimal hyperplane that distinctly categorizes the classes. In binary classification, a hyperplane can be considered as a boundary delineating two distinct data classes. While numerous hyperplanes might achieve this separation, the algorithm selects the one that provides the most effective separation. For a specific classification purpose, as is the case for this project, we have subsequently used SVC, a type of SVM specialized for classifications.
@@
    DT: Each internal node represents a feature or attribute, and each branch represents a decision rule based on that feature. The leaf nodes of the tree represent the final class label or predicted value. During training, each value is separated based on the attribute. When making predictions, new data points traverse the decision tree by following the decision rules at each node until reaching a leaf node, which then provides the predicted class label or value.
@@
    NB classifier: this classifier uses probability to predict whether an input will fit into a certain category. It builds a statistical model based on these probabilities. Naive Bayes calculates the likelihood of the data point belonging to each class using the previously estimated probabilities.
@@
Traditional classifiers have big advantages for small or medium datasets that require simpler or linear models. They have few layers in their architecture; conversely, deep learning (MLP and Transformers) architectures comprise multiple layers of neural networks. Deep architectures take advantage of unsupervised pre-training at the layer level, which facilitates efficient tuning of the deep networks and enables them to extract intricate structures from input data. These extracted features at higher levels contribute to improved predictions and overall performance [36]. For classification, MLP and Transformers are neural networks classifiers:
@@
  • MLP: it consists of multiple layers of nodes (neurons) that are interconnected through weighted connections. MLP employs a feedforward mechanism, where information flows from the input layer through the hidden layers to the output layer. Each node in the network applies an activation function to the weighted sum of its inputs to produce an output. Through a process called backpropagation, the MLP classifier adjusts the weights to minimize the error between predicted and actual labels during training.
@@
  • Transformers: it relies on the attention mechanism. The attention-mechanism looks at an input sequence and decides at each step which other parts of the sequence are important. A Transformer is an architecture for transforming one sequence into another one with the help of two parts (Encoder and Decoder).
@@
The objective is to apply all these classifiers to the PPG signal pulses so that a comparison of the classifiers on our medical data can be built. In addition to being compared with each other, these classifiers will also be compared with the LP semi-supervised algorithm, which annotates our database and classifies artifacts in PPG signals.
IV Experimental Implementation
@@
First, as a reminder, in the LP algorithm, the two input matrices are the annotation matrix, a binary vector, and a matrix containing the features for each pulse. Each pulse represents a node in the algorithm’s graph. For the choice of features, the signal from a temporal perspective has been considered. Therefore, an input matrix for the algorithm of size 256 samples $\times$ the number of pulses can be obtained.
@@
Different metrics have been chosen to evaluate our results. The negative state (0) is a pulse without motion artifacts, whereas the positive state (1) is a pulse with motion artifacts. All these measures are based on the evaluation of false negatives (FN), pulse with artifact incorrectly identified as a clean pulse, false positives (FP), clean pulse incorrectly identified as a pulse with artifact, true negatives (TN), clean pulse correctly identified as a clean pulse, and true positives (TP), pulse with artifact correctly identified as a pulse with artifact. The following metrics are defined:
@@
  • Precision, Recall, and F1: these three scores give a more general idea of how the algorithm works, rather than just looking at the algorithm’s accuracy, which can be biased in certain situations. They are defined as follows:
@@
  • Matthews Correlation Coefficient (MCC), Cohen’s Kappa Coefficient, and Critical Success Index (CSI): MCC is particularly useful in the case of binary classification, where the two classes are unbalanced. It varies between 0 and 1. CSI is also known as the Threat Score. A CSI of 1 indicates perfect prediction, while a score of 0 indicates no successful predictions beyond random chance. Kappa Coefficient ($\kappa$) is stronger than accuracy; it ranges from -1 to 1.
    $\text{mcc}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$
    $\text{csi}=\frac{TP}{TP+FN+FP}$
[…]
@@
  • AUROC/AUC: the AUROC or AUC (Area Under the Receiver Operating Characteristic curve) represents the probability that the model correctly ranks a randomly chosen positive instance higher than a randomly chosen negative instance. The ROC curve is created by plotting the TP rate against the FP rate.
TABLE II: Summary table containing the Label Propagation parameters
Model Kernel Number of neighbors Number of iterations Convergence threshold
Label Propagation KNN 7 1000 $10^{-3}$
@@
Using these metrics, the hyperparameters of our model for the LP algorithm need to be defined. The first parameters defined are the parameters of the function used. The choice of the kernel is between KNN or RBF (radial basis function); depending on this choice, two other associated parameters could be modified. The maximum number of iterations and the algorithm’s convergence tolerance remained the default values: 1000 iterations and $10^{-3}$ for the convergence threshold. The first parameter to be defined was the choice of kernel. For this, a cross-validation was carried out on the data. This involves dividing the data into several parts and then running the two algorithms using different values for the parameters on each part, keeping one part aside for performance testing. Then a calculation of the average performance can be done over all the test parts for each value and choose the one that gives the best performance. A KNN kernel with a number of neighbors of 7 has been chosen. The table II summarizes the parameters chosen for the algorithm. The data are separated as follows: 70% training and 30% testing. The data from the training part are redivided evenly to obtain 50% of unlabeled data and 50 % of labeled data.
V Results and Discussion